# Masked Autoencoders Are Effective Tokenizers for Diffusion Models

### v1

@@ 1 Introduction
Motivated by these insights, we demonstrate that diffusion models trained on AEs with discriminative latent space are enough to achieve SOTA performance. We propose to train AEs as as Masked Autoencoders (MAE) (He et al., 2022; Xie et al., 2022; Wei et al., 2022), a self-supervised paradigm that can discover more generalized and discriminative representations by reconstructing proxy features (Zhang et al., 2022). More specifically, we adopt the transformer architecture of tokenizers (Yu et al., 2021, 2024c; Li et al., 2024c; Chen et al., 2024a) and randomly mask the image tokens at the encoder, whose features need to be reconstructed at the decoder (Assran et al., 2023). To maintain a pixel decoder with high reconstruction fidelity, we adopt auxiliary shallow decoders that predict the features of unseen tokens from seen ones to learn the representations, along with the pixel decoder which is normally trained as previous tokenizers. The auxiliary shallow decoders introduce trivial computation overhead during training. This design allows us to extend the MAE objective that reconstructs masked image patches, to simultaneously predict multiple targets, such as HOG (Dalal & Triggs, 2005) features (Wei et al., 2022), DINOv2 features (Oquab et al., 2023), CLIP embeddings (Radford et al., 2021; Zhai et al., 2023), and Byte-Pair Encoding (BPE) indices with text (Huang et al., 2024).
@@
  • SOTA Generation Performance: Diffusion models of 675M parameters trained on MAETok with 128 tokens achieve performance comparable to previous best models on 256$\times$ 256 ImageNet generation and outperform 2B USiT at 512 resolution with a 1.69 gFID and 304.2 IS.
@@ 2 On the Latent Space and Diffusion Models
To study the relationship of latent space for diffusion models, we start with popular tokenizers, including AE (Hinton & Salakhutdinov, 2006), VAE (Kingma, 2013), representation aligned VAE, i.e., VAVAE (Yao & Wang, 2025). We train diffusion models on them and establish connections between latent space properties and the quality of the final image generation through empirical and theoretical analysis.
Empirical Analysis. Inspired by existing theoretical work (Chen et al., 2022, 2023; Benton et al., 2024), our investigation of the connection between latent space and generation quality starts with a high-level intuition. With optimal diffusion model parameters, such as sufficient total time steps and adequately small discretization steps, the generation quality of diffusion models is dominated by the denoising network’s training loss (Chen et al., 2022, 2023; Benton et al., 2024), while the effectiveness of training diffusion model via DDPM (Ho et al., 2020) heavily depends on the hardness of learning the latent space distribution (Shah et al., 2023; Diakonikolas et al., 2023; Gatmiry et al., 2024). Specially, when the training data distribution is too complex and multi-modal, i.e., not discriminative enough, the denoising network may struggle to capture such entangled global structure of latent space, resulting in degraded generation.
@@
Building upon this intuition, we use the Gaussian Mixture Models (GMM) to evaluate the number of modes in alternative latent space representations, where a higher number of modes indicates a more complex structure. Fig. 2a analyzes the GMM fitting by varying the number of Gaussian components and comparing their negative log-likelihood losses (NLL) across different latent spaces, where a lower NLL indicates better fitting quality. We observe that, to achieve comparable fitting quality, i.e., similar GMM losses, VAVAE requires fewer modes compared to VAE and AE. Fewer modes are sufficient to adequately represent the latent space distributions of VAVAE compared to those of AE and VAE, highlighting simpler global structures in its latent space. Correspondingly, Fig. 2b reports the training losses of diffusion models with AE, VAE, and VAVAE, which (almost) align with the GMM losses shown in Fig. 2a, where fewer modes correspond to lower diffusion losses and better gFID. This alignment validates our intuition, confirming that latent spaces with fewer modes and thus more separated and discriminative features can reduce the learning difficulty and lead to better generation quality of diffusion models.
@@ 4.4 Main Results
Generation. We compare SiT-XL and LightningDiT based on variants of MAETok in Tables 2 and 3 for the 256$\times$256 and 512$\times$512 ImageNet benchmarks, respectively, against other SOTA generative models. Notably, the naive SiT-XL trained on MAETok with only 128 tokens and plain AE architecture achieves consistently better gFID and IS without using CFG: it outperforms REPA (Yu et al., 2024d) by 3.59 gFID on 256 resolution and establishes a SOTA comparable gFID of 2.79 at 512 resolution. When using CFG, SiT-XL achieves a comparable performance with competing autoregressive and diffusion-based baselines trained on VAEs at 256 resolution. It beats the 2B USiT (Chen et al., 2024b) with 256 tokens and also achieves a new SOTA of 1.69 gFID and 304.2 IS at 512 resolution. Better results have been observed with LightningDiT, trained with more advanced tricks (Yao & Wang, 2025), where it outperforms MAR-H of 1B parameters and USiT of 2B parameters without CFG, achieves a 2.56 gFID and 224.5 IS, and 1.72 gFID with CFG. These results demonstrate that the structure of the latent space (see Fig. 4), instead of the variational form of tokenizers, is vital for the diffusion model to learn effectively and efficiently. We show a few generation samples in Fig. 1, with more visualization included in Section C.4.
@@ 4.5 Discussion
as shown by the unconditional generation performance in Table 5. As the latent space becomes more discriminative, the unconditional generation performance also improves significantly. This implies that the CFG linear combination scheme may become less effective, aligning with our CFG tuning results included in Section C.2.
@@ 5 Related Work
Image Generation. The paradigms of image generation mainly categorize to autoregressive and diffusion models. Autoregressive models initially relied on CNN architectures (Van den Oord et al., 2016) and were later augmented with Transformer-based models (Vaswani et al., 2023; Yu et al., 2024b; Lee et al., 2022; Liu et al., 2024; Sun et al., 2024) for improved scalability (Chang et al., 2022; Tian et al., 2024). Diffusion models show strong performance since their debut Sohl-Dickstein et al. (2015b). Key developments (Nichol & Dhariwal, 2021; Dhariwal & Nichol, 2021; Song et al., 2022) refined the denoising process for sharper samples. A pivotal step in performance and efficiency came with latent diffusion (Vahdat et al., 2021; Rombach et al., 2022b), which uses tokenizers to reduce dimension and conduct denoising in a compact latent space (Van Den Oord et al., 2017; Esser et al., 2021; Peebles & Xie, 2023). Recent advances include designing better tokenizers (Chen et al., 2024a; Zha et al., 2024; Yao & Wang, 2025) and combining diffusion with autoregressive models (Li et al., 2024b).
@@ B.2 Training Details of Diffusion Models
We present the training details of SiT-XL and LightningDiT in Tables 7 and 8, which mainly follows their original setup.
@@ C.3 Latent Space Visualization
C.4 More Qualitative Generation Results

### v2

@@ 1 Introduction
Motivated by these insights, we demonstrate that diffusion models trained on AEs with discriminative latent space are enough to achieve SOTA performance. We propose to train AEs as Masked Autoencoders (MAE) (He et al., 2022; Xie et al., 2022; Wei et al., 2022), a self-supervised paradigm that can discover more generalized and discriminative representations by reconstructing proxy features (Zhang et al., 2022). More specifically, we adopt the transformer architecture of tokenizers (Yu et al., 2021, 2024c; Li et al., 2024c; Chen et al., 2024a) and randomly mask the image tokens at the encoder, whose features need to be reconstructed at the decoder (Assran et al., 2023). To maintain a pixel decoder with high reconstruction fidelity, we adopt auxiliary shallow decoders that predict the features of unseen tokens from seen ones to learn the representations, along with the pixel decoder which is normally trained as previous tokenizers. The auxiliary shallow decoders introduce trivial computation overhead during training. This design allows us to extend the MAE objective that reconstructs masked image patches, to simultaneously predict multiple targets, such as HOG (Dalal & Triggs, 2005) features (Wei et al., 2022), DINOv2 features (Oquab et al., 2023), CLIP embeddings (Radford et al., 2021; Zhai et al., 2023), and Byte-Pair Encoding (BPE) indices with text (Huang et al., 2024).
@@
  • SOTA Generation Performance: Diffusion models of 675M parameters trained on MAETok with 128 tokens achieve performance comparable to previous best models on 256 ImageNet generation and outperform previous 2B USiT at 512 resolution with 1.69 gFID and 304.2 IS.
@@ 2 On the Latent Space and Diffusion Models
To study the relationship of latent space for diffusion models, we start with popular tokenizers, including AE (Hinton & Salakhutdinov, 2006), VAE (Kingma, 2013), representation aligned VAE, i.e., VAVAE (Yao & Wang, 2025). We train our own AE and VAE tokenizers under the same training recipe and the same dimension for fair comparison. We train diffusion models on them and establish connections between latent space properties and the quality of the final image generation through empirical and theoretical analysis.
Empirical Analysis. Inspired by existing theoretical work (Chen et al., 2022, 2023; Benton et al., 2024), our investigation of the connection between latent space and generation quality starts with a high-level intuition. With optimal diffusion model parameters, such as sufficient total time steps and adequately small discretization steps, and with assumed similar capacity of tokenizer decoders, the generation quality of diffusion models, i.e., the learned latent distribution, is dominated by the denoising network’s training loss (Chen et al., 2022, 2023; Benton et al., 2024), while the effectiveness of training diffusion model via DDPM (Ho et al., 2020) heavily depends on the hardness of learning the latent space distribution (Shah et al., 2023; Diakonikolas et al., 2023; Gatmiry et al., 2024). Specially, when the training data distribution is too complex and multi-modal, i.e., not discriminative enough, the denoising network may struggle to capture such entangled global structure of latent space, resulting in a degraded generation quality.
@@
Building upon this intuition, we use the Gaussian Mixture Models (GMM) to evaluate the number of modes in alternative latent space representations, where a higher number of modes indicates a more complex structure. The details of GMM training are included in Section B.3. Fig. 2a analyzes the GMM fitting by varying the number of Gaussian components and comparing their negative log-likelihood losses (NLL) across different latent spaces, where a lower NLL indicates better fitting quality. We observe that, to achieve comparable fitting quality, i.e., similar GMM losses, VAVAE requires fewer modes compared to VAE and AE. Fewer modes are sufficient to adequately represent the latent space distributions of VAVAE compared to those of AE and VAE, highlighting simpler global structures in its latent space. Correspondingly, Fig. 2b reports the training losses of diffusion models with AE, VAE, and VAVAE, which (almost) align with the GMM losses shown in Fig. 2a, where fewer modes correspond to lower diffusion losses and better gFID. This alignment validates our intuition, confirming that latent spaces with fewer modes and thus more separated and discriminative features can reduce the learning difficulty and lead to better generation quality of diffusion models.
@@ 4.2 Design Choices of MAETok
We include more ablation study on the number of learnable latent tokens and 2D RoPE in Section C.4.
@@ 4.4 Main Results
Generation. We compare SiT-XL and LightningDiT based on variants of MAETok in Tables 2 and 3 for the 256$\times$256 and 512$\times$512 ImageNet benchmarks, respectively, against other SOTA generative models. Notably, the naive SiT-XL trained on MAETok with only 128 tokens and plain AE architecture achieves consistently better gFID and IS without using CFG: it outperforms REPA (Yu et al., 2024d) by 3.59 gFID on 256 resolution and establishes a SOTA comparable gFID of 2.79 at 512 resolution. When using CFG, SiT-XL achieves a comparable performance with competing autoregressive and diffusion-based baselines trained on VAEs at 256 resolution. It beats the 2B USiT (Chen et al., 2024b) with 256 tokens and also achieves a new SOTA of 1.69 gFID and 304.2 IS at 512 resolution. Better results have been observed with LightningDiT, trained with more advanced tricks (Yao & Wang, 2025), where it outperforms MAR-H of 1B parameters and USiT of 2B parameters without CFG, achieves a 2.56 gFID and 224.5 IS, and 1.72 gFID with CFG. When using a USiT-2B (Chen et al., 2024b) for 512 generation, it pushes the gFID without CFG to 1.72, and gFID with CFG to 1.65. These results demonstrate that the structure of the latent space (see Fig. 4), instead of the variational form of tokenizers, is vital for the diffusion model to learn effectively and efficiently. We show a few selected generation samples in Fig. 1, and more uncurated visualizations are included in Section C.5.
@@ 4.5 Discussion
as shown by the unconditional generation performance in Table 5. As the latent space becomes more discriminative, the unconditional generation performance also improves significantly. This implies that the CFG linear combination scheme may become less effective (Zhao & Schwing, 2025), aligning with our CFG tuning results included in Section C.2. Moreover, adopting more recent advanced CFG techniques, such as Autoguidance (Karras et al., 2024) with naively earlier checkpoints of the generative model, and guidance-free training (Chen et al., 2025) can further improve the gFID of the SiT-XL model from 1.67 to 1.54 and 1.51, respectively. We leave the exploration of other CFG techniques for future work.
Learnable Tokens, Mask Modeling, and Auxiliary Decoders. We also provide a study of the effect of each component in MAETok, as shown in Table 6. The results reveal clear and complementary gains from each design choice. When all three are present, MAETok reaches a better trade-off between the reconstruction and generation quality; dropping the auxiliary decoder or the masking objective, for instance, leads to noticeably weaker results, while omitting the learnable tokens, i.e., using only 256 image tokens, impairs both fidelity metrics most severely.
Mask
Modeling
Learnable
Token
Aux.
Decoder
rFID gFID
0.85 5.78
0.64 8.44
1.01 6.85
1.15 17.18
0.43 9.88
0.96 18.23
0.67 24.47
Table 6: Effects of different components in MAETok.
@@ 5 Related Work
Image Generation. The paradigms of image generation mainly categorize to autoregressive and diffusion models. Autoregressive models initially relied on CNN architectures (Van den Oord et al., 2016) and were later augmented with Transformer-based models (Vaswani et al., 2023; Yu et al., 2024b; Lee et al., 2022; Liu et al., 2024; Sun et al., 2024) for improved scalability (Chang et al., 2022; Tian et al., 2024). Diffusion models show strong performance since their debut Sohl-Dickstein et al. (2015b). Key developments (Nichol & Dhariwal, 2021; Dhariwal & Nichol, 2021; Song et al., 2022) refined the denoising process for sharper samples. A pivotal step in performance and efficiency came with latent diffusion (Vahdat et al., 2021; Rombach et al., 2022b), which uses tokenizers to reduce dimension and conduct denoising in a compact latent space (Van Den Oord et al., 2017; Esser et al., 2021; Peebles & Xie, 2023; Qiu et al., 2025). Recent advances include designing better tokenizers (Chen et al., 2024a; Zha et al., 2024; Yao & Wang, 2025) and combining diffusion with autoregressive models (Li et al., 2024b).
@@ 6 Conclusion
Acknowledge
The authors would like to thank the anonymous reviewers and area chair for their helpful comments. This research project has benefited from the Microsoft Accelerating Foundation Models Research (AFMR) grant program. Difan Zou and Yujin Han acknowledge the support from NSFC 62306252, Hong Kong ECS award 27309624, Guangdong NSF 2024A1515012444, and the central fund from HKU.
@@ B.2 Training Details of Diffusion Models
We present the training details of SiT-XL and LightningDiT in Tables 8 and 9, which mainly follows their original setup.
@@
B.3 Training Details of GMM Models
In Fig. 2, we train our own AE, KL-VAE, and MAETok under exactly the same settings and use the pre-trained VAVAE (Yao & Wang, 2025). The evaluation in Fig. 2 is performed with the same latent size and input dimensions. Specially, for GMM in Fig. 2a, we first represent the original latent size as $(N,H,C)$, where $N$ refers to the training sample size, $H$ refers to the number of tokens, and $C$ refers to the channel size. Following the typical GMM training, we performed the following steps: (1) Latents flatten: The latent size becomes $(N,H\times C)$. (2) Dimensionality Reduction: To avoid the curse of dimensionality, we consider PCA and select a fixed dimension $K$ that results in an explained variance greater than 90%. This step makes the latent dimension $(N,K)$, ensuring that all latent spaces have consistent dimensions. (3) Normalization: To avoid numerical instability and feature scale differences, we further standardize the latent data. (4) Fitting: We fit the data using GMM and return the negative log-likelihood losses (NLL). We train the GMM on the entire Imagenet with a batch size of 256 on a single NVIDIA A8000. It should be noted that distributed training would further optimize the fitting time. The training time for GMM components of 50, 100, and 200 is roughly 3, 8, and 11 hours, respectively.
For SiT-L loss in Fig. 2b, we train SiT-L on the latent space of these four tokenizers for 400K iterations, using an optimizer of AdamW, a constant learning rate of 1e-4, and no weight decay.
@@ C.3 Latent Space Visualization
C.4 More Ablation Results
Tok 1D # Tokens rFID gFID
VAVAE 256 0.28 13.65
MAETok 256 0.37 5.05
MAETok 256 1.01 6.85
MAETok 128 0.48 5.69
(a) Latent tokens.
Pos. Emb. 256 rFID 512 rFID
APE 0.73 1.43
RoPE 0.51 0.72
(b) RoPE.
Table 13: Ablations of latent tokens and 2D RoPE with MAETok on 256$\times$256 ImageNet. We report rFID of tokenizer and gFID of SiT-L trained on latent space of the tokenizer without classifier-free guidance. We train tokenizer of 250K and SiT-L for 400K steps.
We present the ablation study on latent tokens and 2D RoPE in Table 13. One can observe from Table 13a that using learnable latent tokens is more effective than using image tokens only, and 128 latent tokens is enough to achieve similar reconstruction and downstream generation performance, compared to 256 tokens. Furthermore, 2D RoPE helps to generalize better on different resolutions, when trained with mixed resolution images.
C.5 More Qualitative Generation Results